OpenStudy (anonymous):

Solve for x. 20x - 2 = 18x - 4 Select one of the options below as your answer: A. x = -2 B. x = -1 C. x = 1 D. x = 2

OpenStudy (anonymous):

its B

OpenStudy (anonymous):

20x-18x=-4+2 solve it for x

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

does this help you out?

OpenStudy (anonymous):

yahh thanks

OpenStudy (anonymous):

2x=-2 x=-2/2 x=-1

OpenStudy (mathteacher1729):

Out of curiosity, are you allowed to use calcs on the tests where they ask you to solve problems like these? Most TI calcs can do this automatically. (But I still believe it's best to actually learn how to obtain solutions independent of calcs.)

OpenStudy (anonymous):

aye, if you dont learn then when it appers later in the futuer and you don't have access to a comp then your stuck in a rut with no were to turn